OpenStudy (happpily):

what are the solutions of the quadratic equation 2x^2-16x+32=0?

OpenStudy (asett):

Do you have any idea where to start?

OpenStudy (happpily):

honestly no.....

OpenStudy (asett):

Okay, first, we can simplify the expression by removing a 2.

OpenStudy (asett):

We would then have the equation:\[2(x ^{2}-8x+16)\] right?

OpenStudy (asett):

* =0

OpenStudy (asett):

So, now we can split the x^2-8x+16 into the two groups. What do you see that makes a 16 through multiplication and an 8 through addition/subtraction?

OpenStudy (happpily):

WAIT I think i got it !!!!! can you check my work?

OpenStudy (asett):

Of course!

OpenStudy (happpily):

2x2 – 16x + 32 = 0 2(x^2-8x + 16) = 0 (x^2-8x+16) = 0 (x^2-4x -4x + 16) = 0 (x(x-4)-4(x-4)) = 0 (x-4)(x-4) = 0 (x-4)^2 = 0 (x-4)^2 = 0 x - 4 = 0 x = 4

OpenStudy (asett):

Whoo! Good job! It's correct!

OpenStudy (happpily):

YAAAAy thank you! :D

OpenStudy (asett):

Your welcome! Any time.